What a contrast,  last week it was cold, wet and windy, this Sunday was warm and sunny and the bees were out enjoying themselves.  It’s a shame that there were only 7 of us to enjoy it.  The bees were on their best behaviour, and we finished at around 16:30. Thanks again to Pippa for a welcome cup of tea at the end.
Weather,  Hot and Sunny.   no wind.    21 C.
Hive # 1   
Inspected by Liz H
Brood chamber ( 14×12)
Queen found at last and marked Blue
0  queen cells
2 frames of brood ( reducing from 8) all stages but not much young brood.
3 frames in the 14×12
room to lay ok
supers 2  but not enough to extract
Temper 9
entrance reduced with fern to allow guard bees to repel large number of wasps around.
Hive #2
Inspected by Mel
Brood chamber 14×12
Checked 3 supers only.
#1 empty but starting to draw,
#2 drawn and starting to fill
#3 full and put above cleared board with porter escapes. They worked ok and no bees left next day.
Inspected by Richard W
Brood chamber:   Deep
Queen not seen. possibly absent but frame of eggs put in on 7/6 not used to make an emergency queen cell, so possibly a virgin queen still around.  Will use one of the mating Nucs if nothing seen on the next inspection.
No eggs or larva, but still a few sealed brood.
stores ok
1# super  quarter full
Temper:  10
Hive #4
Inspected by Brad D
Brood Chamber: 2 Deeps
Queen seen in bottom deep
2 taken down in bottom deep
4 frames brood
room for laying ok
Super#  1  filling
Temper: 8
Notes: Top deep full of stores, so put above Rhombus clearer board and taken off for extracting next day. clearer board worked ok with only a few bees left.    Entrance also reduced.
Inspected by Mel
Brood chamber #  2 Deep
Queen not seen
no queen cells or brood
plenty of stores and room
Temper : 8
notes: Top deep put under bottom deep
Inspected by Richard W
Brood chamber:  1 Deep
Queen seen and repainted as very faded.
No Queen cells
5 frames of BAS and new eggs
4 frames of stores
2 frames of room
supers : 2  S1 three quarters full,  S2  quarter full.
Notes:  Some (10 +) chalk brood found on brood frames.
Temper: 10
Inspected by Mel
All Super hive. –  only supers checked.
Top super full and put above Clearer board with porter bee escapes.  2nd super  three quarters full.
Temper: 9
Notes: –  Another problem with roof ventilation mesh missing, arrived next day to a robbing mass of bees. cappings and dead bees blocking the porter bee escapes.  Roof replaced and new Canadian clearer board put under.
Inspected by Brad D.
Brood Chamber: Shallow/Deep.  Not Checked
Supers#3   s3 undrawn, s2 and s1 checked and full frames moved to s1 for putting above Canadian clearer board.
Entrance reduced
Notes:  Canadian clearer board worked very well, with very few bees left.
Inspected by Liz H
Ware Hive,
QNS,  but no Queen cells
Brood all stages with day old eggs.
8 frames of stores
plenty of room
Temper : 10
Notes:  2 torn down queen cells located,  lot of wasp activity, entrance reduced by 50%
Nuc  15B
Moved another meter , Deep box, making up nicely.
5 frames Brood all stages and new eggs, stores ok and still room.
Nuc  15F
Housed in Deep,  5 frames of Deep, building up ok BAS and new eggs, but queen not seen
stores and room ok
Temper: 8
Poly Nuc     still going – frames of BAS and new eggs.  QNS
Nuc 15 D  and  Nuc 15 G       Both shaken out as no brood or eggs, so queenless.